OpenStudy (anonymous):

Solve. check for extraneous solutions. √ x+√ 2x=√ 2x

OpenStudy (anonymous):

u'l get extraneous root when you square both sides but if you don't then u'l get no extraneous roots

OpenStudy (anonymous):

can you show it to me @divu.mkr

OpenStudy (anonymous):

x+2x +2x(sqrt2)= 2x

OpenStudy (anonymous):

not really can you help me out lol im bad at math

OpenStudy (mathstudent55):

Is this the problem? \( \sqrt{x} + \sqrt{2x} = \sqrt{2x} \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Then start by subtracting \( \sqrt{2x} \) from both sides. What do you get?

OpenStudy (anonymous):

-2 0 ?

OpenStudy (mathstudent55):

\(\sqrt{x} + \sqrt{2x} = \sqrt{2x}\) \(\sqrt{x} = 0\) Now square both sides, x = 0 There's your solution.

OpenStudy (anonymous):

thanks a lot:)

OpenStudy (mathstudent55):

You're welcome. We didn't square both sides, so we did not introduce any extraneous solutions.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

lol so whats next haha

OpenStudy (mathstudent55):

This problem is done. You can check the solution, x = 0, in the original equation and you'll see that it works.

OpenStudy (anonymous):

ow thanks again

OpenStudy (mathstudent55):

You're welcome.